Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image


Mission Operations Days: 2011/225 through 2011/231
August 13, 2011 00:00:00 GMT to August 19, 2011 00:00:00 GMT

Aqua Spacecraft and MODIS Instrument Status:

PM-1 "Aqua" is in Fine Pointing Normal Mode
FM1 MODIS is in B-Side Nominal Science Mode
BlackbodyA Off; B On (285K)Nominal Science Mode
Calibration ElectronicsA Off; B OnNominal Science Mode
Control ProcessorA Off; B OnNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ClosedNominal Science Mode
FDDI FormatterA Off; B OnNominal Science Mode
FDDI PortA Off; B OnNominal Science Mode
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 Off, #2 OnNominal Science Mode
PV VIS FPAA Off; B OnNominal Science Mode
PV NIR FPAA Off; B OnNominal Science Mode
PV SMIR FPAA Off; B OnNominal Science Mode
PV LWIR FPAA Off; B OnNominal Science Mode
PC LWIR FPAA Off; B OnNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOffNominal Science Mode
RC SMIR FPA HeaterOn (83K)Nominal Science Mode
Scan AssemblyA Off; B OnNominal Science Mode
SDSMOffNominal Science Mode
  • 10W Lamps - 1, 4 (functional)
  • 10W Lamps - 2, 3 (non-functional)
  • 1W Lamps - 1, 2 (functional)
Nominal Science Mode
PS1 Survival HeatersEnabledNominal Science Mode
PS2 Survival HeatersEnabledNominal Science Mode
Timing GeneratorA Off; B OnNominal Science Mode
Flight SoftwareRev BJ + 1 set of patches
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
Nominal Science Mode
Inhibit Ids SetNoneNominal Science Mode

This Week's Completed MODIS Activities:

Saturday, August 13, 2011


Sunday, August 14, 2011

2011/226 21:12:56-21:15:27 MCL MOD_OA27_PV_ECAL.1
OA-27: Photovoltaic bands electronics calibration
Centered on S/C maximum latitude at 21:14:11.

Monday, August 15, 2011

2011/227 20:21:58-20:22:14 MCL MOD_SRCA_SETUP_L1_L4_L5_CC.1
SRCA setup activity using lamps 1, 4, and 5, and in constant current
feedback mode.
2011/227 20:24:58-20:55:23 MCL MOD_OA19_SRCA_FULL_RAD.2
OA-19: SRCA Full Radiometric Calibration - B Side
Data collection 20:40:59-20:55:22 (Centered on S/N Night)
S/C Night 20:31:11-21:05:11.

Tuesday, August 16, 2011


Wednesday, August 17, 2011


Thursday, August 18, 2011


Friday, August 19, 2011


This Week's Scheduled MODIS Activities Not Completed:


This Month's Upcoming MODIS Events:

Calibrations for August will include an SDSM Screened (DOY 220), an SDSM Opened (DOY 220), an SRCA Spectral (DOY 223), a PV ECAL (DOY 226), and an SRCA Radiometric (DOY 227).


There will be an Aqua Drag Make Up Maneuver (DMU) #51 on DOY 237 (8/25/11). The Aqua spacecraft slew out start time is set for 237/14:35:00 and slew back end time is set for 237/15:44:10. The maximum yaw angle for this maneuver will reach 14.35 degrees. The preliminary burn duration for the maneuver is currently 24.25 seconds. The preliminary loss of pointing accuracy times for this maneuver are 237/14:35:00 - 15:59:10.

MODIS Anomalies:


General Instrument Comments:

FM1 MODIS is in Science Mode. The SVD and NAD are open, the survival heaters are enabled, and the Black Body is heated to 285 K.

MODIS Telemetry Trends:

The FM1 MODIS instrument telemetry is trending nominally.

Non-MODIS Significant Events:

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links webpages. There are directions at the URL below for how to obtain a username/password.

Limited Life Item Status:

SRCA 10W Lamp #1: 293.3 of 500 hours
SRCA 10W Lamp #2: 188.0 of 500 hours
SRCA 10W Lamp #3: 205.7 of 500 hours
SRCA 10W Lamp #4: 104.5 of 500 hours

SRCA 1W Lamp #1: 523.4 of 5000 hours
SRCA 1W Lamp #2: 274.9 of 5000 hours

Solar Diffuser Door: 2954 of 3022 Movements
Nadir Aperture Door: 1053 of 1316 Movements
Space View Door: 632 of 1316 Movements